OpenStudy (anonymous):

Will give medal if you help! Two boys on bicycles start from the same place and ride in opposite directions. One rides twice as far as the other one. In 3 hours they are 27 kilometers apart. What is the rate of speed for each boy?

OpenStudy (jojo4eva):

Sorry @Boyce i was wrong

OpenStudy (anonymous):

That is ok @jojo4eva keep trying.:)

OpenStudy (anonymous):

Let the speeds be a, and b. The distances should then be 3a and 3b. So 27 = 3a+3b We also know that a = 2b since one is twice as fast.

OpenStudy (jojo4eva):

lol i think @wio has u sorry i had to fix my calculations

OpenStudy (anonymous):

ok keep going @wio

OpenStudy (anonymous):

You don't know how to solve systems of equations?

OpenStudy (anonymous):

oh yes i do. Give me one sec and I will have it...:)

OpenStudy (anonymous):

Got it @wio Is this right: 27=3x +3y 27=3(2y)+3y 27=6y+3y 27=9y y=3 x=2(3) x=6

OpenStudy (anonymous):

You can always check by putting it back into the equations.

OpenStudy (anonymous):

Sounds good. thanks so much!